trish30:

Anne & Mary went for a walk. Anne started walking 1 minute earlier than Mary did. Anne walks at a speed of 6 feet per second. Mary walks at a speed of 8 feet per second. How many seconds had Anne been walking at the point when the two girls had walked the exact same distance?

trish30:

oh so it would be 10 @Ultrilliam

Ultrilliam:

Just out of curiosity, are there options to the Question?

trish30:

@Ultrilliam no

trish30:

@Ultrilliam so that's all we have to do

Ultrilliam:

Didn't think so, but the answer is 4 seconds, as the math demonstrates, the one minute is a red herring so to speak

Ultrilliam:

If I read the Question correct that is

trish30:

@Ultrilliam I thought the answer was 240

trish30:

because following your way would give Anne walked for x + 60 = 180 + 60 = 240 sec

Ultrilliam:

@Eiwoh2 Sorry I think I might be doing this wrong... Math is not my best subject, I'll leave this to him (the one I tagged) Welcome to QC and sorry I could be more help!

Eiwoh2:

Why u tag me? :\ I am terrible at Math, Ultri XD Ill leave this to @Hero

Ultrilliam:

Hero isn't on...

Eiwoh2:

@Elsa213

Ultrilliam:

@ThisGirlPretty

Eiwoh2:

or @theDeviliscoming :\ idk

theDeviliscoming:

I'm still horrible at math

Eiwoh2:

We apologize for not being able to help you at the moment @trish30 ^^;

trish30:

@Eiwoh2 @Ultrilliam that's ok

Ultrilliam:

You happened to catch us at a brief moment when all our math users are offline, which rarely happens, RIP

trish30:

@elsa213 I don't understand that site's explanation

Eiwoh2:

HERO COME BACK ON!!! ;-;

Ultrilliam:

@vocaloid @Nnesha They should see this next they are on

Elsa213:

`We know that distance = rate x time We know Anne walks at a rate of 6 ft/s and Mary walks at a rate of 8 ft/s 1 min = 60 sec, and Anne started walking 60 sec prior to Mary If the amount of time Mary walked is x, then the amount of time that Anne walked is (x + 60) Since we want to know when the two girls walked the same distance, we can form the following equation using the formula for distance 8x = 6(x + 60) 8x = 6x + 360 2x = 360 x = 180 So Mary walked for 180 sec and Anne walked for x + 60 = 180 + 60 = 240 sec`

Ultrilliam:

So it was 240 o.o

trish30:

@elsa213 thank you so much

Eiwoh2:

Elsa = Flawless

Elsa213:

Ofc. :D Welcome to QuestionCove. c:

Hero:

I should have known @Elsa213 did not figure this out on her own. SMH

Elsa213:

Eraser >.>

Hero:

Pencil >.>

theDeviliscoming:

Chalk >.>

Elsa213:

Board >.>

theDeviliscoming:

Chocolate >.>

Hero:

See what you've started @Elsa213

Elsa213:

Luff o;

theDeviliscoming:

Unicorns

Elsa213:

I didn't start it though. o: You started and I continued. e.e

Hero:

It started with you copy and pasting answers from some other site.

Elsa213:

It started with you having a life and not able to show up when you were tagged.

Hero:

SMH

Elsa213:

cx <3